OpenStudy (anonymous):

A theater has 24 rows of seats. There are 16 seats in the first row and each successive row increases by 2 seats, 1 on each side. a) calculate the number of seats in the 24th row. b) calculate the number of seats in the whole theater. Please show work!

Directrix (directrix):

The attached file shows the formulas you need. http://www.regentsprep.org/Regents/math/algtrig/ATP2/ArithSeq.htm

Directrix (directrix):

a) calculate the number of seats in the 24th row. a sub 24 = 16 + (24 - 1)*2 = 62 seats n the 24 the row b) S sub 24 = 24 ( 16 + 24) / 2 = 12* 40 = 480 seats total Check my work, please.

OpenStudy (anonymous):

Sorry for getting back to you so late, everything looks good!

Directrix (directrix):

Thanks.